OpenStudy (anonymous):

3x2 +5x = 6 Can someone please solve this problem and show all the work? Please use factoring or the quadratic formula. I am having a hard time getting the correct answer so i need to compare my work with someone else's

OpenStudy (anonymous):

3x² + 5x - 6 = 0 3x² + 6x - x - 6 = 0 solve further to get the factors/roots

OpenStudy (anonymous):

if u need more help, pls inform...

OpenStudy (anonymous):

3x² + 5x - 6 = 0 3x² + 6x - x - 6 = 0 solve further to get the factors/roots

OpenStudy (anonymous):

-5 +- root(72) all over 6 will give aprox...

OpenStudy (anonymous):

if I didn't make a mistake.. a = 3 , b = 5 , c = -6

OpenStudy (anonymous):

use the quadratic eq.

OpenStudy (anonymous):

3x² + 5x - 6 = 0 3x² + 6x - x - 6 = 0 3x(x + 2) -1(x + 6) = 0 (3x - 1)(x + 2) = 0 if 3x -1 = 0, the x = 1/3 if x + 2 = 0, then x = -2 so the two roots are 1/3 and -2

hero (hero):

Wait...what happend to the x+6 ?

OpenStudy (anonymous):

Sorry, made a big boo-boo.....

OpenStudy (anonymous):

Correct solution is... 3x² + 5x - 6 = 0 here a = 3 , b = 5 and c = -6 - b +/- root(b² - 4ac) - 5 +/- root(25 - 4*3*(-6) x = -------------------- = ------------------------ 2a 2*3 -5 +/- root(25 + 72) -5 +/- root(97) x = -------------------- = ----------------- 6 6